### v2

# reference.bib

Abstract
State-of-the-art large language models (LLMs) are commonly deployed as online services, necessitating users to transmit informative prompts to cloud servers, thus engendering substantial privacy concerns. In response, we present ConfusionPrompt, a novel private LLM inference framework designed to obfuscate the server by: (i) decomposing the prompt into sub-prompts, and (ii) generating pseudo prompts along with the genuine sub-prompts as input to the online LLM. Eventually, the returned responses can be recomposed by the user to obtain the final whole response. Such designs endows our framework with advantages over previous protocols that (i) it can be seamlessly integrated with existing black-box LLMs, and (ii) it achieves significantly better privacy-utility trade-off than existing text perturbation-based methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirement for a privacy-preserving group of prompts, and provide a complexity analysis, affirming ConfusionPrompt’s efficiency. Our empirical evaluation reveals that our method offers significantly higher utility compared to local inference methods using open-source models and perturbation-based techniques, while also requiring much less memory than open-source LLMs.
1 Introduction
Large language models (LLMs) have shown immense potential in various tasks [brown2020language, roziere2023code, huang2023instruct2act]. As a result, they are progressively being incorporated into various real-world applications, including medical consultation [thirunavukarasu2023large] and financial service [wu2023bloomberggpt]. Currently, state-of-the-art LLMs such as ChatGPT from OpenAI [ouyang2022training] are typically provided as online services, since LLM owners deserve the intellectual property of their models and thus they are unwilling to reveal model parameters to the clients [openai2023gpt4]. However, direct usage of such online service could raise significant privacy concerns, since the users’ prompts are likely to hold sensitive information that should remain concealed from the server.
Available solutions to privacy-preserving model inference include encryption [liu2023llms, chen2022x] and perturbation [du2023dp] techniques, which are impractical for private LLM inference. In encryption methods, the server leverages homomorphic encryption (HE) [acar2018survey] and Secure Multiparty Computation (MPC) [cramer2015secure] to conduct private inference on the users’ encrypted query. The practicality of this method is limited by its high computation overheads, especially on LLMs (e.g., GPT-3 has 175B parameters [brown2020language]). The basic idea of a perturbation-based method is to inject a specific level of noise to user’s input before releasing it to the server, which is challenging to strike a nice privacy-utility balance [lyu2020differentially].
Another issue with the above solutions is that they rely on the service provider to modify their framework or release specific model parameters, which could be extremely costly. The application of encryption techniques relies on platform’s investment in hardware and algorithm acceleration. Perturbation-based methods often rely on local differential privacy (LDP) [dwork2006differential] to ensure privacy with formal guarantees. However, recent studies have proposed deploying certain modules on the user side for improved LDP performance, necessitating the sharing of specific model parameters by the server [qu2021natural, mai2023split, malekzadeh2024salted]. These requirements impose additional burdens and proprietary concerns on the service provider, hindering the deployment of these solutions.
To alleviate the above concerns, this paper proposes a novel private LLM inference framework ConfusionPrompt that can be seamlessly integrated with existing online black-box LLMs. The basic idea of ConfusionPrompt is to construct a group of prompts (containing real and fake prompts) with privacy guarantee to confuse the server, making the curious server fail to infer the real user prompt. However, a key challenge arises when a prompt contains multiple private attributes. In such cases, the user needs to consider all combinations of true and fake attributes to effectively confuse an attacker with prior knowledge of certain attributes, rendering the query complexity exponential in the attribute size. To address this issue, we propose a decomposition approach in which the query is divided into sub-questions. This ensures that the private information is distributed across different sub-queries, thereby reducing the complexity to approximately linear in the attribute size (see Figure 4).
Following the above ideas, our proposed ConfusionPrompt consists of four critical steps: (1) The user decomposes the original prompt into several sub-prompts; (2) For each genuine sub-prompt, the user generates a series of pseudo prompts that obfuscate the sensitive attributes in the genuine sub-prompt, until meeting the requirement of privacy guarantee. (3) User queries online services with a group of prompts, receives the corresponding responses, and retrieves the sub-responses for the genuine sub-prompts. (4) User recompose the sub-responses to obtain the final response. Such a framework empowers the user with independent deployment, eliminating the need for additional privacy-preserving actions from the service provider.
Accordingly, we develop a $(\lambda,\mu,\rho)$-privacy model for ConfusionPrompt to formulate the requirements for a privacy-preserving group of prompts. Through complexity analysis, we show that for a prompt $\boldsymbol{p}$ with $U(\boldsymbol{p})$ private attributes and $\mu$ privacy budget, the basic confusion strategy without decomposition requires to generate $\mathcal{O}\left((1/\mu)^{U(\boldsymbol{p})}\right)$ fake prompts. While our decomposition strategy in ConfusionPrompt can reduce the complexity to $\mathcal{O}\left((1/\mu)U(\boldsymbol{p})\right)$ in the ideal decomposition scenario. Overall, based on the privacy model and complexity analysis, we derive the criteria for an ideal decomposer and generator, and accordingly construct a preference dataset to train them using direct preference optimization (DPO) [rafailov2023direct].
Our contributions are as follows:
(1) We are the first to propose a private LLM inference framework using confusion-based strategy. Our ConfusionPrompt framework can be seamlessly implemented by clients with existing online black-box LLMs such as ChatGPT and Claude with improved privacy-utility trade-off.
(2) To safeguard user’s privacy, we define a privacy model to formulate the requirements of a group of prompts.
(3) We leverage a local decomposition module to reduce the complexity of the prompt group under the same privacy protection level.
(4) Experiments show that our ConfusionPrompt can achieve consistently and significantly better utility than local inference method using open-source models and LDP-based methods. Additionally, ConfusionPrompt requires significantly lower memory cost compared to open-source LLMs.
2 Related Works
In this section, we review three types of methods for privacy-preserving LLM inference: anonymization-based method, encryption-based method, and perturbation-based method.
Anonymization-based method: Traditional anonymization techniques rely on Named Entity Recognition (NER) to redact Personally Identifiable Information (PII) such as individuals’ names, social security number, and email addresses [ehrmann2023named, lample2016neural]. Recent research on LLM inference involves sanitizing sensitive items in the input and subsequently de-anonymize the LLM’s returned responses [kan2023protecting, chen2023hide]. The downside of this method is that anonymization fall short in concealing other sensitive information, including verbs and non-named entities [anandan2012t].
Encryption-based method: Cryptonets [gilad2016cryptonets] proposed the first neural network inference on encrypted data using homomorphic encryption (HE). They approximated the non-linear function such as Sigmoid and MaxPooling by polynomials. Iron [hao2022iron] designed specialized and efficient protocols for two types of computationally heavy operations in Transformer-based inference: (i) matrix multiplications, and (ii) complex functions including Softmax, GELU activations, and LayerNorm. To achieve further speedup, [liu2023llms] transformed the high-overhead functions into cryptography-friendly approximations, and finetune the model to maintain accuracy.
Perturbation-based method: Perturbation-based method provide privacy guarantee by inject calibrated noise into the input. Existing research utilizing this method predominantly focus on privacy protection in fine-tuning [kerrigan2020differentially, yu2021differentially] and prompt-tuning phases [duan2023flocks, li2023privacypreserving], while few research investigate the privacy-preserving inference paradigm. A major challenge in perturbation-based private inference is to balance the utility and privacy tradeoff. Recent studies have proposed Text2Text [feyisetan2020privacy] and paraphraser-based approaches [utpala2023locally, mattern2022limits] to privatize text with LDP. Split-N-Denoise (SnD) [mai2023split] deployed the token embedding layer at the client side, and introduced a user-side denoising model to correct the purturbed embedding output for downstream tasks.
3 ConfusionPrompt: Design and Formulation
In this section, we first illustrate the overview of our proposed ConfusionPrompt, which consists of three critical components: decomposer, generator, and recomposer. Then, we develop a privacy model to formulate the requirements for achieving privacy-preserving inference.
3.1 Overall Design
Denote $G_{s}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ as the large LM deployed as cloud service such as ChatGPT [ouyang2022training, openai2023gpt4], where $\mathcal{V}^{*}$ is the vocabulary space. Instead of sending raw user prompt to the cloud, our ConfusionPrompt introduces a new privacy-preserving inference paradigm, which makes it significantly hard for any third party to infer the real original user prompt even if it happens to access the information sent from the user. Basically, ConfusionPrompt involves the following six steps:
(1) Decomposition of the original prompt. We first introduce a decomposer $G_{d}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ that aims to decompose the user’s original prompt $p$ to a sequence of genuine sub-prompts $\boldsymbol{p}=[p_{1},p_{2},...,p_{|\boldsymbol{p}|}]$, resulting in fewer private attributes (i.e., private information) in each sub-prompt.
(2) Generation of pseudo prompts. A pseudo prompt generator $G_{f}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ is introduced to generate a pseudo prompt given a genuine sub-prompt by replacing specified critical information. By generating multiple pseudo prompts for each genuine sub-prompt and mixing them together as a prompt group, we are able to hide the genuine sub-prompts, increasing the difficulty of being inferred. For the simplicity of privacy analysis, we design to enforce the pseudo prompt to be consistent with the genuine prompt in terms of syntactic structure.
(3) Evaluation of privacy level. To ensure that the genuine prompt is safely hidden in the prompt group, we also need to design some criteria (e.g., semantic irrelevance) to evaluate the usability of generated pseudo prompts. Based on these criteria, we will continue sample and filter the pseudo prompts until the prompt group meet the privacy guarantee, which will be introduced in details in 3.2.
(4) Communication with cloud server. Here, the user sends the prompt group to the server for the usage of large LM on the cloud, while the server subsequently sends back the response group to the user. During this process, we expect that it will be extremely hard for the server to infer the original user prompt according to the prompt group, therefore safeguarding user’s privacy.
(5) Retrieval of interested response. Since the user itself is aware of which sub-prompts are genuine, it can seamlessly retrieve the corresponding sub-responses and discard the responses of pseudo prompts.
(6) Recomposition of sub-responses. Here, we introduce a recomposer $G_{r}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ that maps a sequence of sub-prompt-response pairs to a final response.
Figure 1: Overview of ConfusionPrompt. (1) Decomposition of the original prompt into sub-prompts. (2) Generation of pseudo prompts. (3) Querying online LLMs to get the sub-responses and retrieving the interested sub-responses. (4) Recomposition of sub-responses to obtain the final response.
3.2 Privacy Model
Here, we define a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements that the group of pseudo prompts should satisfy for privacy protection, which offers guidance for the evaluation of privacy level and training of models (i.e., decomposer, generator, and recomposer) (i) The fabricated attributes in the pseudo prompt should be semantically irrelevant to the original attributes in genuine prompt, so that the genuine user information can be effectively obfuscated [wu2015constructing]. (ii) The genuine prompt should be obfuscated by sufficient pseudo prompts, and the attacker can not easily identify the real prompt from the combination pattern even with background knowledge. (iii) The pseudo prompt should appear genuine rather than obviously fabricated. Thus it’s crucial to maintain the fluency and reasonability of the pseudo prompts.
In the following, we first give the definitions of private attributes and their semantic relevance.
Definition 3.1 (Private Attributes).
Denote $\mathcal{U}$ as the attribute space, and $\mathcal{P}$ as the prompt space. Given a sub-prompt list $\boldsymbol{p}=\in\mathcal{P}$, the contained private attributes are denoted as $U(\boldsymbol{p})=\{u_{1},u_{2},...,u_{m}\}$ for $m$ attributes.
Remark 3.2.
For conciseness, the prompt in definition 3.1 could be a full prompt $p$, or a sequence of decomposed prompts with the same intention $\boldsymbol{p}=[p_{1},p_{2},...,p_{|\boldsymbol{p}|}]$.
The private attributes encompass a range of elements, including verbs, adjectives, and nouns, which extend beyond personally identifiable information (PIIs). Next, we define the semantic similarity between two private attributes.
Definition 3.3 (Attribute-attribute Similarity).
Given two attributes $u_{1}$, $u_{2}\in\mathcal{U}$, an attribute similarity function measures the semantic similarity between $u_{1}$ and $u_{2}$: $\rm{Sim}(u_{1},u_{2}):\mathcal{U}\times\mathcal{U}\rightarrow\mathbb{R}$.
For a pair of genuine and pseudo prompts, we define the correspondent attributes as followed:
Definition 3.4 (Correspondent Attributes).
Given two prompts $\boldsymbol{p}_{1}$, $\boldsymbol{p}_{2}\in\mathcal{P}$, the correspondent attribute of $u_{i}\in U(\boldsymbol{p}_{1})$ in $\boldsymbol{p}_{2}$, is defined as the attribute at the same syntactic position of $\boldsymbol{p}_{2}$, denoted by $\rm{Corr}(u_{i},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$.
Remark 3.5.
The definition assumes that: (1) $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ have the same syntactic structure; (2) if $u_{i}$ occurs in $\boldsymbol{p}_{1}$ multiple times, the correspondent locations in $\boldsymbol{p}_{2}$ return the same attribute.
For example, given the original question ‘What are the responsibilities of software engineers?’ and private attribute ‘software engineer’, the correspondent attribute in the fake question ‘What are the responsibilities of school teachers?’ would be ‘school teacher’. Based on the correspondent attributes and attribute-attribute similarity, we can define the semantic similarity between two prompts below:
Definition 3.6 (Prompt-prompt Similarity).
Given two prompts $\boldsymbol{p}_{1}$, $\boldsymbol{p}_{2}\in\mathcal{P}$, their similarity can be defined by the similarity between each pair of correspondence attributes:
$\mathrm{Sim}(\boldsymbol{p}_{1},\boldsymbol{p}_{2})=\max_{u_{i}\in U(\boldsymbol{p}_{1})}\rm{Sim}\left(u_{i},\rm{Corr}(u_{i},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\right)$ (1)
We proceed to the definition of significance of an attribute, which is tied to the likelihood of a curious server being able to identify the genuine attribute from a group of genuine and pseudo prompts.
Definition 3.7 (Significance of Single Attribute).
Denote =$\mathbb{P}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{n}\}$ as a group of prompts. The significance of an attribute $u\in U(\boldsymbol{p}_{i})$ related to a group of prompts $\mathbb{P}$ is defined as:
$\mathrm{Sig}(u,\mathbb{P})=\frac{\sum_{j=1}^{n}H(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j})}{n},$ (2)
where:
$H(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j})=\left\{\begin{array}[]{ll}1&\rm{Corr}\left(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j}\right)=u\\ 0&\rm{otherwise}\end{array}.\right.$ (3)
In our setting, the group of prompts consists of one genuine prompt along with a collection of pseudo prompts. The significance of a true attribute can be considered as the proportion of its occurrence within the group of prompts. In the following, we provide the definition of significance for multiple attributes.
Definition 3.8 (Significance of Attribute Set).
Suppose $\boldsymbol{p}_{0}$ denotes the genuine prompt. The significance of attribute set $U(\boldsymbol{p}_{i})$ related to a group of prompts =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ is defined as:
$\mathrm{Sig}(U(\boldsymbol{p}_{0}),\mathbb{P})$ (4)
$=$ $\max_{u_{k}}\max_{h}\max_{\mathcal{V}}\frac{\sum_{j=1}^{n}\bar{H}(\mathcal{V},p_{0h},p_{jh})}{\sum_{j=1}^{n}\bar{H}(\mathcal{V}\backslash u_{k},p_{0h},p_{jh})}$
${\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0h}),$
where:
$\bar{H}(\mathcal{V},p_{0h},p_{jh})=\left\{\begin{array}[]{ll}1&\mathrm{Corr}\left(u,p_{0h},p_{jh}\right)=u\ \forall u\in\mathcal{V}\\ &\mathrm{or}\ \mathcal{V}=\emptyset\\ 0&\mathrm{otherwise}\end{array},\right.$ (5)
$p_{ih}$ is the $h$-th decomposed sub-prompt in the $i$-th prompt. In case that the attacker have prior knowledge of some private attributes, we introduce a notation $\mathcal{V}$ as a subset of private attributes in $p_{0h}$. Then we consider the proportion of a given attribute conditioning on any possible set of private attributes to bound the possibility of correctly identify the target attribute with some background knowledge. Now, we provide the definition of genuineness for the third criterion:
Definition 3.9 (Genuineness).
Denote $D:\mathcal{P}\rightarrow\mathbb{R}$ as a function to discriminate between fabricated and genuine prompts. A larger value returned by $D$ indicates a higher likelihood of a prompt being genuine. The genuineness of a prompt $p\in\mathcal{P}$ can be defined as $\rm{Genu}(p)=D(p)$.
Based on definition 3.6, 3.8, and 3.9, we can formulate our proposed $(\lambda,\mu,\rho)$-privacy model as:
Definition 3.10 (User Privacy).
Let =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are pseudo prompts. If $\mathbb{P}$ satisfies the following requirements, then it is deemed that it can ensure $(\lambda,\mu,\rho)$-privacy of user prompt $\boldsymbol{p}_{0}$ with respect to discriminator $D$:
  • Each pseudo prompt should be semantically irrelavant to user prompt $\boldsymbol{p}_{0}$, i.e., $\forall i\in[1,n]$, $\rm{Sim}(\boldsymbol{p}_{0},\boldsymbol{p}_{i})\leq\lambda$.
  • The users’ sensitive attributes should be obfuscated by sufficient pseudo prompts, i.e., $\rm{Sig}(U(\boldsymbol{p}_{0}),\mathbb{P})\leq\mu$.
  • Each pseudo prompt should not be classified as fabricated prompt by discriminator $D$, i.e., $\forall i\in[1,n]$, $\rm{Genu}(p_{i})=D(p_{i})\geq\rho$.
In Appendix C, we provide the relationship between our privacy model and the upper bound on the success rate of inference attacks, which states that an curious server can do no better than random guessing if the combination of attributes are uniformly distributed.
4 ConfusionPrompt: User-Side Models
Following the overall design and the requirements in privacy model, this section introduces how to enable ConfusionPrompt from the perspective of designing user-side models, including decomposer, generator, and recomposer.
4.1 Decomposer
A decomposer is designed to decompose an original user prompt into several sub-prompts, such that the required number of pseudo prompts to ensure the same level of privacy preservation can be significantly reduced. In the following, we theoretically demonstrate the benefit of a local decompose module in terms of complexity measured by the number of required pseudo prompts.
4.1.1 Complexity Analysis
We first show the complexity for single-paragraph prompt, that is, the prompt is not decomposed.
Theorem 4.1 (Complexity for Single Paragraph).
Let =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a single paragraph $|\boldsymbol{p}_{i}|=1$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires that
$n\geq\left(\frac{1}{\mu}\right)^{|U(\boldsymbol{p}_{0})|}.$ (6)
Accordingly, we provide the following complexity analysis for decomposed prompts, which are a key component in our ConfusionPrompt.
Theorem 4.2 (Complexity for Decomposed Prompt).
Let =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a sequence of decomposed prompt with the same intention $\boldsymbol{p}_{i}=[p_{i1},p_{i2},...,p_{il}]$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires that
$n\geq\sum_{j=1}^{l}\left(\frac{1}{\mu}\right)^{|U(p_{0j})|}$ (7)
Specifically, if $|U(\boldsymbol{p}_{0})|=l$ and $|U(p_{0j})|=1$, $\forall j\in[1,l]$, then
$n\geq\frac{|U(\boldsymbol{p}_{0})|}{\mu}$ (8)
Theorem 4.1 and 4.2 compares the complexities between single prompt and decomposed sub-prompt, where the proof is defered to appendix B. In the ideal scenario, the complexity can be reduced to $\mathcal{O}\left(|U(\boldsymbol{p}_{0})|/\mu\right)$ after decomposition. See Figure 4 for intuitive explanation.
4.1.2 Requirements of Ideal Decomposer
Now, we can articulate the criteria for an ideal decomposer:
(1) A fundamental requirement for the decomposition process is to follow the collectively exhaustive principle [rasiel1999mckinsey]. In other words, the collection of responses to decomposed prompt should recover the whole response.
(2) An ideal decomposer would optimize the complexity, i.e., the required number of pseudo prompts, in two aspects: (i) each sub-prompt contain as few attributes as possible; (ii) each attribute appears in as few prompts as possible.
4.1.3 Model Training
Following the spirit of Reinforcement Learning from Human/AI Feedback (RLHF/RLAIF) [christiano2017deep, lee2023rlaif], we train a decomposer that optimizes the aforementioned two objectives simultaneously. For training efficiency and stability, the reward model training and reinforcement learning stages are replaced with Direct Preference Optimization (DPO), which directly optimizes a model on the preference data [rafailov2023direct]. Therefore, the training prosedure of decomposer is as follows:
Supervised fine-tuning (SFT): We finetune a pretrained language model (LM) on demonstration data for decomposition task, where the LM is trained to decompose given a full prompt.
Preference data construction: We use the SFT model to generate various decompositions given a full prompt $p$ to generate a various decompositions $(p_{1},p_{2},...,p_{l})\sim\pi_{\mathrm{SFT}}(\cdot|p)$. The comparison data considering the two requirements/dimensions are collected by either human labelers or AI feedback, following the rules in section D.1.
DPO optimization: Given a full prompt $p$, we denote $p_{w}$ and $p_{d}$ as the preferred and dispreferred decomposition. Then the optimization objective is formulated as:
$\begin{gathered}\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta};\pi_{\mathrm{SFT}})=-\mathbb{E}_{(p,p_{w},p_{d})}\\ \left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(p_{w}|p)}{\pi_{\mathrm{SFT}}(p_{w}|p)}-\beta\log\frac{\pi_{\theta}(p_{d}|p)}{\pi_{\mathrm{SFT}}(p_{d}|p)}\right)\right]\end{gathered}$ (9)
where $\sigma$ is the logistic function, $\beta$ is a hyperparameter controlling the deviation from SFT model, and $\theta$ is the model parameters to optimize.
4.2 Generator
Generator takes an genuine prompt $p$ and its private attributes $U(p)$ as input and produces a pseudo prompt $p^{\prime}$ with replaced attributes. The generator will be run for multiple times to obtain a series of pseudo prompts for each genuine prompt.
Requirements. The requirements for a pseudo prompt suggest the following criteria for an ideal generator: (1) The generator should produce pseudo prompt with fake information that’s semantically irrelevant to the correspondent attributes in genuine prompt. (2) The pseudo prompt should not be classified as fabricated prompt by a strong discriminator $D$.
Model training. The generator is trained using a similar procedure described in Section 4.1.3: (i) SFT: fine-tune a pretrained LM using demonstration data for replacement task, where the LM is trained to generate new prompt with replaced attributes. (ii) Preference data construction: collect comparison data considering both semantic relevance and genuineness; see details in section D.2). (iii) DPO optimization: optimize the model on the comparison data.
4.3 Recomposer
The local re-composition LM combines the retrived sub-prompts and sub-responses to produce the final whole response. We train a local recomposer on a collection of demonstration data with supervise learning. Specifically, given a complete prompt $p$ and a set of sub-prompt-response pairs $\left((p_{1},r_{1}),...(p_{k},r_{k})\right)$, the recomposer $G_{r}$ is trained to maximize $\log p_{G_{r}}\left(r|p,(p_{1},r_{1}),...(p_{k},r_{k})\right)$ where $r$ is the final response.
4.4 Algorithm
Algorithm 1 outlines the procedure for our private inference framework ConfusionPrompt. The client starts with decomposing the full prompt into sub-prompts, and generate a collection of qualified fake prompts that meets the privacy requirement. It can be derived immediately from the algorithm that our protocol guarantees $(\lambda,\mu,\rho)$-privacy.
5 Experiment
5.1 Experiment Setup
We evaluate our framework on two multi-hop reasoning datasets: StrategyQA [geva2021did] and MuSiQue [trivedi2022musique]. We consider ConfusionPrompt with different online LLMs including GPT-4, GPT-4-Turbo, and GPT-4o [openai2023gpt4]. We consider the privacy budget $\lambda\in[0.5,0.8]$, $1/\mu\in[5,50]$, and $\rho\in[1,4]$ unless specified. Refer to section F.2 for the training of decomposer, generator, and recomposer.
5.2 Empirical Privacy Evaluation
We simulate two types of inference attacks to investigate the privacy protection level under various combinations of privacy parameters.
Prompt Identification Attack: an attack that identifies the true query from a group of fake and true prompts. Denote $\mathbf{q}=[\![q_{1}||q_{2}||...||q_{n}]\!]$ as the concatenated group of queries, and $k$ as the index for the true query. We finetune a BART-large classification model $G_{\rm{PIA}}$ to predict the index of the true query $\hat{k}=G_{\rm{PIA}}(\mathbf{q})$.
Attribute Inference Attack: an attack that infers the sensitive features of records from either a group of prompts, or differentially privatized inputs. We rely on the twitter text dataset [vashisth2020gender] to predict the gender based on the user’s review.
Noted that prompt identification attack is specifically designed for our framework, and attribute inference attack can be applied to both ConfusionPrompt and DP-based methods.
5.3 Experiment Results
5.3.1 Privacy Experiments
Figure 2 visualizes the attack accuracy for prompt identification attack under various combinations of significance $\mu$, genuineness $\rho$, similarity $\lambda$. It can be observed that: (1) decreasing significance $\mu$ consistently reduces the attack accuracy. (1) As the genuineness threshold $\rho$ increases from 1 to 4, the attack accuracy approaches that of random guessing. The result reveals that it becomes harder for the attacker to distinguish the true query as the fake prompts become more fluent. (3) There’s no clear relationship between the attack accuracy and similarity threshold $\lambda$. Such privacy parameter could be more related to the attribute inference attack discussed later.
Figure 2: Prompt identification attack accuracy under various combinations of privacy parameters.
The results of attribute inference attack are detailed in F.4. A key observation is that when $1/\mu\geq 15$, the attack accuracies for ConfusionPrompt decrease to a plateau and is similar to that of Text2Text with $\epsilon\leq 1$.
5.3.2 Utility Evaluation
For StrategyQA, we report the accuracy scores (ACC) and area under the roc curve (AUC) to assess the the classification task. For MuSiQue, we report the F1 score, ROUGE-L, and exact matching (EM) to evaluate the question answering task [yang2018hotpotqa]. We benchmark our framework with $\mu=15$, $\rho=4$, and $\lambda=0.5$ against five baseline methods: (1) a local Llama2 [touvron2023llama] model with 7B parameters; (2) a local Vicuna [chiang2023vicuna] model with 13B parameters; (3) Text-to-text privatization (Text2Text) [feyisetan2020privacy], where the query is privatized with LDP by replacing each word with the perturbed token embeddings; (4) Paraphraser [utpala2023locally, mattern2022limits] that paraphrases the prompt with LDP mechanism using a language model; (5) querying online GPTs with no privacy protection. In the former two methods, the user downloads the open-source models and directly extract the final response from the model. In methods (3) and (4), the queries are privatized with LDP privacy budget $\epsilon=10$ before transmission to the server. For the final approach, the complete prompt is transmitted to the service provider without any confusion strategy.
Table 1: Utility comparisons between our proposed ConfusionPrompt and baselines. ConfusionPrompt strikes the best balance between privacy preservation and utility.
Method Privacy StrategyQA MuSiQue
Protection ACC AUC F1 roughL EM
Llama2-7B Yes 0.602 0.581 0.484 0.485 0.321
Vicuna-13B Yes 0.646 0.628 0.417 0.416 0.225
Text2Text (GPT-3.5-Turbo) Yes 0.528 0.496 0.019 0.019 0.010
Text2Text (GPT-4-Turbo) Yes 0.537 0.506 0.028 0.027 0.014
Text2Text (GPT-4o) Yes 0.533 0.500 0.021 0.020 0.012
Paraphraser (GPT-3.5-Turbo) Yes 0.489 0.478 0.076 0.076 0.028
Paraphraser (GPT-4-Turbo) Yes 0.546 0.523 0.060 0.061 0.026
Paraphraser (GPT-4o) Yes 0.537 0.515 0.072 0.072 0.038
GPT-3.5-Turbo No 0.751 0.737 0.604 0.605 0.439
GPT-4-Turbo No 0.803 0.798 0.663 0.664 0.503
GPT-4o No 0.791 0.784 0.721 0.721 0.557
ConfusionPrompt (GPT-3.5-Turbo) Yes 0.723 0.726 0.606 0.605 0.445
ConfusionPrompt (GPT-4-Turbo) Yes 0.741 0.743 0.633 0.634 0.495
ConfusionPrompt (GPT-4o) Yes 0.733 0.739 0.685 0.684 0.535
The results are reported in Table 1. From the results, we see that (1) online proprietary models outperform open-source models by over 11% and 24%, repectively, for StrategyQA and MuSiQue, indicating the strong motivation for users to use online service while preserving privacy. (2) ConfusionPrompt based on various online models consistently and significantly outperforms LDP-based methods, showing the effectiveness of our proposed method. Notably, our framework possesses a distinct advantage over perturbation-based methods: increasing the privacy budget does not adversely affect performance since the true queries are always provided to the server.
5.3.3 Decomposition Benefits
Theorem 4.1 and 4.2 show that the decomposition module could reduce the number of pseudo prompts required to be sent to the server. In this section, we empirically demonstrate practical benefits by analyzing the monetary cost of our ConfusionPrompt framework. In particular, we compute the monetary ratio, the ratio of the monetary cost for using ConfusionPrompt compared to the direct query method, both before and after the application of the decomposition module.
Figure 3 presents the query cost benefits achieved through our decomposition module. It can be observed that the complexity gains become increasingly pronounced as the level of privacy protection increases. This can be attributed to the fact that the complexity experiences polynomial growth before decomposition, whereas it exhibits approximate linearity after decomposition. Our protocol offer substantial savings ranging from an average reduction of 1.9×1.9$\times$1.9 × at $\mu=1/10$ to an average reduction of 18.3×18.3$\times$18.3 × at $\mu=1/50$.
Figure 3: Monetary ratio of strategyQA and MuSiQue dataset before (decomp) and after (w/o decomp) decompositon under various privacy budget $\mu$. Decompostion in ConfusionPrompt significantly reduces the monetary cost, indicating its efficiency.
We evaluate generation on longer attributes, memory cost of local models, the performance of the discriminator $D$ that verifies the genuineness, and pseudo prompt generation cost (see Appendix F.5, F.7, F.8, and F.9).
6 Conclusion
This paper proposes a private inference framework over online LLMs, termed ConfusionPrompt. We deploy three local models on the user side: (i) decomposer that maps an original prompt to a sequence of sub-prompts, (ii) generator that produces pseudo prompts by replacing private attributes in the genuine sub-prompts, (iii) recomposer that maps the decomposed prompt-response pairs from the cloud service to the final response. Such designs endows our framework with advantages over previous protocols that: (i) it can be seamlessly integrated with existing black-box LLMs while the service providers have no need to modify their already-built framework, and (ii) it achieves better privacy-utility trade-off than existing privatization methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirement for a privacy-preserving group of prompts, and accordingly provide a complexity analysis to demonstrate the benefits of the decomposition module. Experiments show that our ConfusionPrompt significantly outperforms the inference with local open-sourced LLMs in terms of utility and computation cost.
[…]
Appendix A Decomposition Example
Figure 4 presents an example demonstrating the decomposition benefits.
Figure 4: Example of decomposition savings in query complexity. Decomposition module reduces the query complexity from 9 to 6 under privacy budget $\mu=3$.
Appendix B Proof of Theorem 4.1 and 4.2
We begin with the proof for Theorem 4.1 as followed:
Proof.
For the prompt with single paragraph, the significance can by represented as:
$\begin{gathered}\mathrm{Sig}(U(p_{0}),\mathbb{P})=\max_{u_{k}}\max_{\mathcal{V}}\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0},p_{i})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\backslash u_{k},p_{0},p_{i})}\\ {\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0}))\end{gathered}$ (10)
, where $p_{0}$ is the genuine prompt, and $p_{i}$ is the $i^{\mathrm{th}}$ prompt in the prompt group, and:
$\bar{H}(\mathcal{V},p_{0},p_{i})=\left\{\begin{array}[]{ll}1&\mathrm{Corr}\left(u,p_{0},p_{i}\right)=u\ \forall u\in\mathcal{V}\\ &\mathrm{or}\ \mathcal{V}=\emptyset\\ 0&\rm{otherwise}\end{array}\right.$ (11)
For $\mathrm{Sig}(U(p_{0}),\mathbb{P})=\mu$, it’s obvious that there should be at least $1/\mu$ distinct values, including genuine and fake values, assigned to each attributes. Otherwise suppose $u_{k}$ have less than $1/\mu$ distinct values, we would have:
$\frac{\sum_{j=1}^{|\mathbb{P}|}H(u_{k},p_{0},p_{j})}{|\mathbb{P}|}>\mu$ (12)
which violates the expression 10.
Next, we claim that each possible combination of attribute set values must appear at least once within the prompt group. To proof by contradiction, let’s suppose that there is one combination missing in the group while the others appear once. Then we can consider two cases:
(a) The missing combination doesn’t contain any true values. Then the significance of other single attribute must be larger than $\mu$. This is because for a set containing all combinations, the proportion of the occurrence for each attribute is exactly $\mu$. Therefore, such removal would make their significance smaller.
(b) The missing combination contain a subset of true values $\mathcal{V}\subseteq U(p_{0})$. Then conditioning on the subset $\mathcal{V}$, the proportion of any other single attribute would be over $\mu$, as their original proportion is $\mu$. Therefore, for any $u_{k}\notin\mathcal{V}$, it holds that:
$\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\cup u_{k},p_{0},p_{i})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0},p_{i})}>\mu$ (13)
which violates the expression 10.
To make each combination occur at least once, we should have at least $(1/\mu)^{|U(p_{0})|}$ prompts in the group. ∎
Following that, we can proceed to the proof for Theorem 4.2:
Proof.
Let $\mu=\mathrm{Sig}(U(p_{0}),\mathbb{P})$ denotes the significance. According to Theorem 4.1, for each sub-query $p_{0h}\in\boldsymbol{p}_{0}$, we should have at least $(1/\mu)^{\|U(p_{0h})\|}$ prompts to satisfy:
$\max_{u_{k}}\\ \max_{\mathcal{V}}\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0h},p_{ih})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\backslash u_{k},p_{0h},p_{ih})}\leq\mu\quad\quad{\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0h})$ (14)
Then the proof completes by summing up the number of prompts over all sub-queries. ∎
Appendix C Bound on Inference Attack
In this section, we consider the attacker identify the target attribute by prior knowledge of other attributes and the combination pattern of attribute values. The following theorem states that an curious server can do no better than random guessing if the combination of values are uniformly distributed:
Lemma C.1.
Suppose the attacker is guessing the attributes from attribute combination pattern in $\mathbb{P}$. The attacker can do no more better than random guessing if the following condition is satisfied:
  • Each combination of attribute set has the same occurrence times in the prompt groups $\mathbb{P}$, i.e., the combinations are uniformly distributed.
From the proof of Theorem 4.1 and 4.2, we can construct a uniformly distributed prompt group under the optimal number of prompts. Therefore, we can turn to the success rate under an random guessing attack, which is bounded by Lemma C.2.
Lemma C.2.
Given a group of prompts =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ with $(\lambda,\mu,\rho)$-privacy, suppose that the attacker have prior knowledge about the attribute subset $\mathcal{V}\subseteq U(\boldsymbol{p}_{0})$. By random guessing, the probability of correctly identifying any target attribute $u\in U(\boldsymbol{p}_{0})\land u\notin\mathcal{V}$ is upper bounded by $\mu$.
Proof.
According to Definition 3.8, it holds that:
$\begin{gathered}\frac{\sum_{j=1}^{n}\bar{H}(\mathcal{V},p_{0h},p_{jh})}{\sum_{j=1}^{n}\bar{H}(\mathcal{V}\backslash u,p_{0h},p_{jh})}\leq\mu\end{gathered}$ (15)
for any $h$ and $u$. In other words, the proportion of any attribute $u$ conditioning on the prior knowledge attribute set $\mathcal{V}$ is upper bounded by $\mu$. Therefore, by random guessing, the probability of correctly identifying any target attribute $u$ is upper bounded by $\mu$. ∎
Appendix D Ranking Rules in Preference Sampling
D.1 Decomposer
We rank the decompositions with the following steps:
  • Identify whether or not each decomposition satisfies the MECE principle.
  • Compute the complexity of each decompositon under a given $\mu$.
  • Rank the responses using the following rules: (i) Decompositions that adhere to the MECEprinciple should be ranked higher than those that do not follow the principle. (ii) If two decompositions both follow or do not follow the MECE principle, the one with lower complexity is preferred.
D.2 Generator
To rank the pseudo prompts, we take the steps below:
  • Compute $s$, the similarity between the pseudo and original prompt.
  • Compute $f$, the genuineness score for the pseudo prompt.
  • Calculate the overall score $\beta f-s$, where $\beta$ is the hyperparameter controlling the weight of the genuiness score.
  • Rank the responses with the overall score, with a higher score indicating a higher preference.
Appendix E Algorithm
In this section we present the algorithm of ConfusionPrompt in 1.
Algorithm 1 ConfusionPrompt
1:  Input: Original prompt $p$, privacy parameters $\lambda$, $\mu$, and $\rho$.
2:  Output: Final response $r$.
3:  User decomposes the original prompt into sub-tasks $\boldsymbol{p}_{0}=G_{d}(p)$.
4:  for $h\in[|\boldsymbol{p}_{0}|]$ do
5:     User keeps sampling from the generator until there are $n$ distinct pseudo prompts with $n$ given by eq. 7, where each one $1\leq i\leq n$ has prompt similarity $\mathrm{Sim}(p_{0h},p_{ih})\leq\lambda$ and genuineness score $\mathrm{Genu}(p_{ih})\geq\rho$.
6:  end for
7:  User sends the group of genuine and pseudo prompts =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ to the server.
8:  Server returns a collection of responses $(r_{0},...,r_{l})$.
9:  User retrieves the response corresponding to the genuine prompt $r_{0}$.
10:  User obtains the final response $r$ using the local recomposition model.
Appendix F Experiment
F.1 Construction of Private Attributes
We extract the private attributes of StrategyQA and MuSiQue through the steps below:
  • Step 1: sample construction. We adopt a semi-automated method to extract private attributes from each sample. Firstly, we extract the entities using a combination of two NER methods: Spacy [srinivasa2018natural] and Flair [akbik2019flair]. Then we manually correct and supplement the private attributes for each question.
  • Step 2: model finetuning. An LLM is finetuned using the private attribute samples to guide the LLM output private attributes given a query.
  • Step 3: attribute extraction. The remaining queries are fed into the finetuned LLM to generate private attributes.
During step 1, we label the private attributes according to the following guidelines:
  • Identify the key components for a query, including but not limited to proper nouns, phrases, verbs, and adjectives.
  • Ensure there is no overlapping information between different private attributes within the same query.
  • Each private attribute should be as succinct as possible. For example, instead of labeling "spouse of the Green performer" as a single attribute, it is preferable to split it into "spouse" and "Green performer" as separate attributes.
The third point is to ensure privacy guarantee even when an attacker possesses prior knowledge of certain attributes. Table 2 provides several examples of private attributes.
Table 2: Examples of private attributes in StrategyQA and MuSiQue.
Query Private Attributes
Are blue lips normal? blue lips, normal
Could ten gallons of seawater crush a six year old? seawater crush, ten gallons, six year old
Who is the spouse of the Green performer? spouse, Green performer
What instrument is played by the person from The Blackout All-Stars? play instrument, The Blackout All-Stars
What is the capital of the county that Pine Springs is located in? capital of the county, Pine Springs
In step 2, we evaluate the generative quality of various LLMs on five models: i) GPT-4-Turbo with few-shot examples, ii) GPT-3.5-Turbo finetuned with 100 samples, iii) BART-large [lewis2020bart], iv) T5-large [raffel2020exploring], and v) Flan-T5-Xlarge [chung2024scaling]. The latter three open-source models are finetuned with 1000 samples.
Table 3 presents the performance on a human-labeled test dataset. It can be observed that the finetuned version of GPT-3.5-turbo gives the best result, and thus it’s used as the attribute extraction model throughout our experiment.
Table 3: Comparison of various attribute extraction models. Finetuning GPT-3.5-turbo returns the highest accuracy.
Method StrategyQA MuSiQue
F1 roughL EM F1 roughL EM
GPT-4-Turbo 0.689 0.748 0.080 0.811 0.762 0.234
GPT-3.5-Turbo 0.803 0.881 0.421 0.846 0.789 0.333
BART-large 0.713 0.690 0.227 0.664 0.634 0.118
T5-large 0.603 0.595 0.102 0.381 0.383 0.010
Flan-T5-large 0.669 0.683 0.243 0.380 0.383 0.026
F.2 Training of local models
F.2.1 Training of decomposer
The decomposition model is finetuned on a pretrained BART model [lewis2020bart] with 406M parameters. During SFT stage, we utilize the decomposition samples from StrategyQA and MuSiQue as demonstration data. Following that, we obtain preference data through the following steps: (i) extract the sub-answers using gpt-4-turbo and obtain the final answers using a recomposition model to judge whether the decomposition returns a correct answer; (ii) compute the complexity measured by the number of required pseudo prompts under $\mu=10$.
During step (i), to address the recomposer’s limitations in determining the correctness of the final answer, we train a robust recomposition model by fine-tuning LLaMa2-7B. Note that during inference, we adopt a much smaller recomposition model for efficiency.
F.2.2 Training of generator
The generator is finetuned on the BART-large model. To collect demonstration data for SFT, we prompt GPT-4-turbo to generate multiple pseudo prompts given the raw prompt and its private attributes using the following template.
Please replace the phrases {attributes} in the each of the following sentences, such that each sentence is fluent and reasonable, and the alterntive phrases have irrelevant meaning as {attributes}.
Please return {# of replaces} replacements for each sentence.
Strictly respond in the form of JSON with the following format: {"S1": ["replacement 1", "replacement 2",…], "S2": ["replacement 1", "replacement 2",…]}.
Sentences: {dictionary of sentences}
For collecting preference data, we employ two models to evaluate the semantic similarity and genuineness, which involves a local similarity evaluation model and discriminator. (1) For similarity evaluation model: We adopt a fine-tuned version of MiniLM-6L model [reimers2019sentence] to extract the embedding of each private attribute. The semantic relevance between a pair of attributes is given by the cosine similarity between their embeddings. (2) For discriminator, We leverage GPT-4-turbo to construct the training dataset for training a local discriminator. As the genuineness is closely related to the sentence’s fluency, we instruct GPT-4-turbo to evaluate the fluency for each sentence [kanumolu2023unsupervised] (see detailed instruction in section F.3).
F.2.3 Training of recomposer
The recomposition model is finetuned on a pretrained RoBERTa [liu2019roberta] and BART-large [lewis2020bart] model, respectively, for strategyQA and MuSiQue. We utilizes the sub-prompts/responses given in StrategyQA and MuSiQue dataset as well as the final response to train a recomposition model. To improve the performance of recomposor, we pretrain the model on two addition datasets, SQuAD [rajpurkar2016squad] and DROP [dua2019drop].
F.3 Semantic Similarity Model and Discriminator
The comparison data collection for the generator involves a local similarity evaluation model and discriminator.
Similarity evaluation model: We adopt a finetuned version of MiniLM-6L model [reimers2019sentence] to extract the embedding of each private attribute. The semantic relevance between a pair of attributes is given by the cosine similarity between their embeddings.
Discriminator: We leverage GPT-4 to construct the training dataset for training a local discriminator. As the genuineness is closely related to the sentence’s fluency, we prompt GPT-4 to evaluate the fluency for each sentence with the following template [kanumolu2023unsupervised]:
Given multiple sentences, use the scoring rules below to score each sentence’s fluency on a scale of 1 to 4:
1. Score 1: Incomprehensible. Inarticulate/ non-fluent sentence.
2. Score 2: Low Quality. Partially fluent sentence: (a) only half of the sentence is fluent or (b) more than 1 missing words or (c) more than 1 misspelt words or d) contains individual fluent word-groups with missing coherence between them.
3. Score 3: Moderate. Sentence is predominantly fluent but contains either (a) misspelt word or (b) missing word or (c) multiple occurrence of a word.
4. Score 4: Perfect. Perfectly fluent sentence without any syntactic or grammatical error.
Strictly respond in the form of JSON with the following format: {"S1": the score, "S2": the score}.
Sentences: {dictionary of sentences}
On obtaining 4000 training and 700 validation samples, we finetune a Bert-base (110M parameters) to train a local discriminator.
F.4 Results on Attribute Inference Attack
Figure 5 presents the attack accuracies of attribute inference attack for both ConfusionPrompt and LDP-based methods, from which we can make the following observations: (1) Paraphrase generally leads to higher attack accuracies compared to Text2text. This can be attributed to the fact that the semantic meanings of certain key words remain unchanged during the paraphrasing process. (2) The attack accuracies decrease to a plateau when $1/\mu\geq 15$ or $\epsilon\leq 1$. This suggests that beyond these thresholds, further increasing the privacy protection measures does not significantly reduce the attack accuracy. (3) The attack accuracy for ConfusionPrompt with $1/\mu\geq 15$ is similar to that of Text2Text with $\epsilon\leq 1$, indicating certain alignment for the privacy protection between LDP-based methods and our framework. (4) Decreasing the similarity threhold $\lambda$ helps to reduce the attribute inference attack accuracies.
Figure 5: Attribute inference attack accuracy for ConfusionPrompt and LDP-based methods.
F.5 Generator for Longer Attributes
F.5.1 Dataset Construction
To explore the generator’s ability on generating longer attributes, we craft a privacy dataset for ConfusionPrompt (PrivCP) with sensitive information from 7 categories. For each category, we utilized GPT-4 to generate 50 simulated user queries, manually inspecting and selecting 20 for necessary grammatical and semantic adjustments. Table 4 presents an example from each category and its associated risks.
Table 4: Sample Questions and Associated Risks for Each Category of PrivCP.
Type N Example Question Potential Risks
Business 20 ‘As a VP of marketing at MovieMagic, I am considering a shift in our advertising spend from TV to digital platforms. What are the key factors to consider?’ Reveal the strategic considerations of the company
Legal 20 ‘I have a meeting with a client at Clark Legal Services in Boston who wants to create a trust. What are the different types of trusts available?’ Reveal the sensitive case details of a client
Health 20 ‘My 3-year-old daughter, Emily, has just been diagnosed with a peanut allergy. How can I ensure she avoids peanuts at her daycare?’ Reveal the medical and personal info of family members
Career 20 ‘I have worked as a civil engineer at AECOM for five years in Los Angeles. What certifications should I pursue for career advancement?’ Reveal the personal career experience and work location
Education 20 ‘I am a parent evaluating private schools in Chicago for my 6-year-old son, Matthew. What are the key factors to consider in choosing a school with special needs support?’ Reveal the personal info and
Social 20 ‘My company is hosting a corporate event at the Hilton in Las Vegas. How can I network effectively with potential clients?’ Reveal the location and purpose of the private business event
Personal 20 ‘I am moving to a penthouse in Manhattan in July. How can I best use natural light in my home decor?’ Reveal the residential location, moving schedule and personal preference for decor
In table 5, we summarize the statistic of private attribute for strategyQA, MuSiQue, and PrivCP in terms of three metrics: i) average number of private attributes per query, ii) average length of each attribute (measured by the number of words), and iii) average attribute-to-query ratio (measured by the percentage of attribute words in a query). It can be observed that PrivCP has much longer attributes than the two standard benchmarks.
Table 5: Statistics of private attributes on three datasets. Avg. stands for average. PrivCP has longer attributes than the two standard datasets.
Avg. # of attributes Avg. length of attributes Avg. attribute- to-query Ratio
strategyQA 2.18 1.74 40.88%
MuSiQue 2.82 2.72 43.20%
PrivCP 2.52 6.75 73.80%
F.5.2 Generation Performance
The generation quality is evaluated with the following measures: i) BLEU [papineni2002bleu] between the generated psuedo sentence and a set of referenced replacement, ii) fluency score on a scale from 1 to 4, and iii) semantic similarity scores ranging from 0 to 1. Details on the computation of fluency and similarity scores are provided in Appendix F.3.
We evaluate the generation performance on three language models in Table 6. It can be observed that while there’s no significant variation of similarity score among models, Flan-T5-large has the best performance in terms of BLEU and fluency score.
Table 6: Generation performance and parameter size of various language models.
BLEU Fluency Score Similarity Score Paratemetr Size
BART-large 0.345 2.9 0.53 406M
T5-large 0.406 3.66 0.557 737M
Flan-T5-large 0.428 3.7 0.555 783M
F.6 Local Model Selection
In this section we justify the selection of decomposor and recomposor. For decomposor, we finetune each candidate model with SFT, and obtain the BLEU between the generated decomposition and reference provided by the dataset. For recomposor, the models are evaluated in terms of the accuracy metrics described in Section 5.3.2, using the golden decomposition and sub-answer as training and testing inputs.
Table 7 presents the experiment results on MuSiQue using models from three different classes of varying sizes. While Flan-T5-large exhibits the best performance for both tasks, we select BART-large as the decomposition and recomposition model due to its balanced trade-off between model size and performance.
Table 7: Parameter size and performance on both decomposition and recomposition tasks across various models for MuSiQue.
Parameter size Decomposer Recomposer
BLEU F1 RougeL EM
BART-base 139M 0.314 0.563 0.564 0.446
BART-large 406M 0.402 0.640 0.641 0.525
T5-base 223M 0.343 0.581 0.580 0.484
T5-large 737M 0.408 0.646 0.645 0.554
Flan-T5-base 248M 0.367 0.608 0.609 0.518
Flan-T5-large 783M 0.458 0.671 0.671 0.577
F.7 Computation Overhead Analysis
To validate the practicality of our method, we investigate the memory cost required for the three local models, as well as that for local open-source LLMs. The results are presented in Table 8. From the results, we see that our proposed ConfusionPrompt is significantly efficient compared to inferring large open-source language models on the user side. Specifically, it requires only $61.2\%$ and $79.7\%$ memory cost compared to running LLaMa2-7B and Vicuna 13B, respectively, on the user side.
Table 8: Memory cost for local inference with 10 samples. Results show that ConfusionPrompt require much less computation overhead compared to inferring large open-source language models on the user side.
Local Model Llama2-7B Vicuna-13B Decomposer Generator Recomposer Total (ConfusionPrompt)
Memory Cost 25885MB 49653MB 3867MB 3865MB 2306MB 10038MB
F.8 Accuracy of Discriminator
In this section, we assess the reliability of the local discriminator with two metric: accuracy (ACC) and mean squared error (MSE). The discriminator is trained on 4000 samples and tested on 700 samples. To validate the capability of a GPT-4 scorer, we constructe 100 samples with manual labeling and obtain the performance of GPT-4 judger on this dataset. According to Table 9, the ACC of local discriminator is above 0.7 for a 4-label classification problem, with MSE approximately 16% of the total scale (0-3). The scores given by GPT-4 judger are also close to the human labels.
Table 9: Accuracy and MSE for local discriminator and GPT-4 judger. The local discriminator is tested on the validation data scored by GPT-4. GPT-4 Judger is evaluated against ground-truth data with manual label.
Local discriminator GPT-4 judger
ACC 0.710 0.800
MSE 0.488 0.200
F.9 Pseudo Prompt Generation Cost
Algorithm 1 suggests that the generator keeps generating fake prompts until the $(\lambda,\mu,\rho)$-privacy is satisfied. Therefore, it’s crucial to consider the number of sampling times it takes for a generator to produce sufficient qualified pseudo prompts.
Table 10: Cost of generation in terms of number of samplings and time cost per sampling (seconds).
$1/\mu$ $5$ $10$ $20$ $30$
Number of Samplings $17.7$ $30.1$ $65.3$ $93.5$
Time per Sampling 1.73 seconds
Table 10 presents the number of generations to perform and the time cost for each generation under different levels of $\mu$. It requires to conduct sampling for an average of $3.23$ times the $1/\mu$. Note that $1/\mu$ can be treated as the minimal number of sampling times to achieve the privacy budget when the prompt contains only one private attribute. The average time per sampling is 1.73 seconds, and thus the overall time consumption for each sample ranges from 30 seconds at $\mu=1/5$ to 161 seconds at $\mu=1/30$.
Appendix G Discussion and Future Work
Extension to diverse task: our framework is experimented on the classification and question-answering task. It’s of significant interest to extend our framework to a broad range of NLP tasks, such as paraphrasing, translation, and summarization. A key challenge in this extension lies in designing an appropriate decomposition methodology for diverse text inputs.
Confusion with historical conversation: we currently consider the one-round conversation setting, while in reality the user may engage in multi-round communications with the server. In such cases, a curious attacker could potentially infer private attributes by detecting patterns in the conversation. It is crucial to develop a confusion strategy along with conducting privacy analysis for multi-round conversations.