### v5

# ConfusionPrompt: Practical Private Inference for Online Large Language Models

Abstract
State-of-the-art large language models (LLM) are often deployed as online services, posing privacy risks for user prompts. In response, we introduce ConfusionPrompt, a novel framework for private LLM inference that protects user privacy by: (i) decomposing the original prompt into smaller sub-prompts, and (ii) generating pseudo-prompts alongside the genuine sub-prompts. ConfusionPrompt seamlessly integrates with existing black-box LLMs and achieves a superior privacy-utility trade-off compared to existing text perturbation methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements for a privacy-preserving group of prompts and provide a complexity analysis to justify the role of prompt decomposition. Our empirical evaluation shows that ConfusionPrompt outperforms local inference with open-source models by over 11% and surpasses perturbation-based techniques by over 42%, while reducing memory consumption by at least 62% compared to open-source LLMs.
{IEEEImpStatement}
The rapid adoption of LLMs via cloud-based services has heightened privacy concerns, as users must transmit sensitive prompts to external servers. Existing approaches for privacy-preserving LLM inference, such as encryption-based techniques and perturbation-based strategies, are often impractical as they either significantly degrade model performance or lack compatibility with black-box LLMs. Our proposed framework, ConfusionPrompt, introduces a novel, client-side framework that enables private inference without requiring any modifications to existing black-box LLM services. ConfusionPrompt consistently outperforms local inference with open-source LLMs and perturbation-based privacy methods in terms of utility, while also minimizing memory and computational overhead. This framework advances toward safer, privacy-conscious deployment of LLMs, an increasingly pivotal concern in real-world AI applications.
{IEEEkeywords}
Large language model, privacy-preserving computation, private inference, confusion strategy
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks [6, 55, 28, 7, 68], driving their adoption in real-world applications such as medical consultations and financial services [67, 66, 59]. State-of-the-art LLMs, such as OpenAI’s ChatGPT [45], are predominantly offered as online services, primarily to safeguard proprietary model parameters and intellectual property [44]. However, this paradigm introduces significant privacy risks, as users often transmit prompts containing sensitive information that should ideally remain confidential from the server.
Existing solutions to privacy-preserving model inference, such as encryption [37, 9] and perturbation [15] techniques, are often impractical for private LLM inference. In encryption-based methods, the server leverages homomorphic encryption (HE) [1] and Secure Multiparty Computation (SMPC) [14] to conduct private inference on the users’ encrypted query. The practicality of this method is limited by its high computation overheads, especially on LLMs (e.g., GPT-3 has 175B parameters [6]). On the other hand, the basic idea of perturbation-based methods is to inject a specific level of noise into the user’s input before releasing it to the server, which is challenging to strike a satisfied privacy-utility balance [40].
Another limitation of the aforementioned solutions is their dependence on the service provider to modify their infrastructure or disclose specific model parameters, which can be prohibitively costly. Encryption-based techniques require significant investment from platform in hardware and algorithmic acceleration to support efficient computation. Perturbation-based methods often rely on local differential privacy (LDP) [18] to ensure privacy with formal guarantees. For improved LDP performance, recent studies have proposed deploying certain modules on the user side, necessitating the sharing of specific model parameters by the server [50, 41, 42, 34]. These requirements impose additional burdens and proprietary concerns on the service provider, hindering the deployment of these solutions. It remains an open challenge to develop a method that seamlessly integrates with existing black-box LLMs while maintaining strong utility.
To alleviate the above concerns, this paper proposes ConfusionPrompt, a novel private LLM inference framework that can be seamlessly integrated with existing online black-box LLMs. The core idea of ConfusionPrompt is to construct a set of prompts, containing real and fake prompts, designed to confuse the server, preventing the server from accurately identifying the genuine user prompt. However, a significant challenge arises when a prompt contains multiple sensitive attributes. In such cases, the user needs to consider all combinations of true and fake attributes to effectively confuse an attacker with prior knowledge of certain attributes, rendering the query complexity growing exponentially with the number of attributes. To address this issue, we propose a decomposition approach in which the query is divided into sub-questions. This ensures that the private information is distributed across different sub-queries, thereby reducing the complexity to approximately linear in the number of attributes (see Figure 2).
Building on the aforementioned concepts, our proposed ConfusionPrompt consists of four critical steps: (1) The user decomposes the original prompt into several sub-prompts. (2) For each genuine sub-prompt, the user generates a set of pseudo prompts that obfuscate sensitive attributes in the genuine sub-prompt, repeating this process until the desired level of privacy is achieved. (3) The user sends a group of both genuine and pseudo-prompts to the online service, retrieves the responses, and isolates the sub-responses corresponding to the genuine sub-prompts. (4) Finally, user recomposes the sub-responses to obtain the final result. This framework allows for independent user-side deployment, eliminating the need for privacy-preserving modifications by the service provider.
Accordingly, we develop a $(\lambda,\mu,\rho)$-privacy model for ConfusionPrompt to formulate the requirements for a privacy-preserving group of prompts. Through complexity analysis, we show that for a prompt $\boldsymbol{p}$ with $U(\boldsymbol{p})$ private attributes and a privacy budget $\mu$, the basic confusion strategy without decomposition requires generating $\mathcal{O}\left((1/\mu)^{U(\boldsymbol{p})}\right)$ fake prompts, while our decomposition strategy in ConfusionPrompt can reduce the complexity to $\mathcal{O}\left((1/\mu)U(\boldsymbol{p})\right)$ in the ideal decomposition scenario. Based on the privacy model and complexity analysis, we derive the criteria for an ideal decomposer and generator, and accordingly design a two-stage training strategy.
Our key contributions are as follows:
(1) We are the first to propose a private LLM inference framework using confusion-based strategy. Our ConfusionPrompt framework can be seamlessly implemented by clients with existing online black-box LLMs, such as ChatGPT and Claude, providing an improved privacy-utility trade-off. In addition, our framework can be integrated with Trusted Execution Environment (TEE) [56] solutions. Even in the event of a TEE compromise, user privacy remains safeguarded through the confusion strategy.
(2) To ensure user privacy, we define a privacy model to formulate the requirements for a group of prompts, including both real and pseudo-prompts. Furthermore, we introduce a local decomposition module to reduce the complexity of the prompt group under the same privacy parameters.
(3) Experiments show that ConfusionPrompt can achieve a consistently and significantly better utility than LDP-based methods. Moreover, ConfusionPrompt surpasses the performance of local inference method using open-source models with lower memory requirement.
2 Related Works
In this section, we review three types of methods for privacy-preserving LLM inference: sanitization-based method, encryption-based method, and perturbation-based method.
Sanitization-based method. Traditional sanitization techniques rely on Named Entity Recognition (NER) to redact Personally Identifiable Information (PII) such as individuals’ names, social security number, and email addresses [19, 32, 48]. The downside of this method is that NER falls short in concealing other sensitive information, including verbs and non-named entities [3]. Recent research on LLM inference involves sanitizing sensitive items in the input and subsequently de-anonymize the LLM’s returned responses [29, 10, 49]. However, such input perturbations can introduce semantic biases, potentially leading to responses that are semantically irrelevant. Casper utilizes a combination of rule-based and machine learning (ML)-based techniques to filter sensitive PIIs and alert users of the sensitive topics [12]. While this method effectively detects sensitive information, it does not address how to process such information without adversely affecting the output of LLMs. In general, sanitization approaches identify and remove sensitive information, but this can degrade output utility when such information is crucial for LLM inference.
Encryption-based method. Cryptonets [23] proposed the first neural network inference on encrypted data using homomorphic encryption (HE). They approximated the non-linear function such as Sigmoid and MaxPooling by polynomials. Iron [24] designed specialized and efficient protocols for two types of computationally heavy operations in Transformer-based inference: (i) matrix multiplications, and (ii) complex functions including Softmax, GELU activations, and LayerNorm. To achieve further speedup, [37] transformed the high-overhead functions into cryptography-friendly approximations, and finetuned the model to maintain accuracy. SecFormer designs a combination of SMPC protocols for accurate and efficient computation of complex nonlinear functions, such as GeLU and LayerNorm [39]. BOLT reduces the payload on both linear and non-linear operations through cryptographic optimizations, and eliminates insignificant words using ML techniques for enhanced efficiency [47]. NEXUS introduces the first non-interactive protocol, allowing the client to complete the entire inference process with the server in a single round of communication [72]. Despite their privacy advantages, encryption-based methods impose significant computational overhead and require substantial investment in hardware and algorithmic acceleration, making integration with existing black-box LLMs challenging.
Perturbation-based method. Perturbation-based methods provide privacy guarantee by injecting calibrated noise into the input. Existing studies utilizing this method predominantly focus on privacy protection in fine-tuning [31, 71, 35] and prompt-tuning phases [17, 36], while few studies investigate the privacy-preserving inference paradigm. A major challenge in perturbation-based private inference is to balance the utility and privacy trade-off. Recent studies have proposed Text2Text [21] and paraphraser-based approaches [64, 43] to privatize text with LDP. Split-N-Denoise (SnD) [41] deployed the token embedding layer at the client side, and introduced a user-side denoising model to correct the purturbed embedding output for downstream tasks. More recently, DYNTEXT proposes a semantic-aware dynamic text sanitization mechanism for privacy-preserving LLM inference, which adaptively perturbs sensitive tokens according to their semantic density to improve the privacy-utility trade-off [73]. InferDPT further studies privacy-preserving inference for closed-box LLMs by combining a perturbation module with a local extraction module, such that a perturbed prompt is uploaded to the remote LLM and a lightweight local model reconstructs a higher-quality output from the perturbed generation [61]. In a related but distinct setting, FedCoT proposes a federated chain-of-thought distillation framework that transfers knowledge from a server-side LLM to a client-side small language model using perturbed prompts and rationales to protect user privacy [20]. However, existing perturbation-based methods either significantly degrade utility or are incompatible with black-box LLMs due to proprietary and efficiency concerns.
In summary, existing approaches either significantly degrade model performance or lack compatibility with black-box LLMs. To address this, we propose a confusion-based strategy that uploads a mix of genuine and synthetic prompts while preserving privacy. Table 1 summarizes the coarse-grained comparision between ConfusionPrompt and existing approaches.
Table 1: Coarse-grained comparison between privacy-preserving LLM inference frameworks.
Method Plug-and-Play Impact on Utility
HaS [10] Yes Medium
Casper [12] Yes Medium
Iron [24] No Low
SecFormer [39] No Low
BOLT [47] No Low
NEXUS [72] No Low
Text2Text [21] Yes High
Paraphraser [64, 43] Yes High
DYNTEXT [73] Yes Medium
InferDPT [61] Yes Medium
SnD [41] No Medium
ConfusionPrompt Yes Low
  ConfusionPrompt offers a plug-and-play solution with little impact on utility. It achieves superior privacy-utility trade-off compared to sanitization and perturbation methods, which often introduce semantic bias into the original prompt.
3 ConfusionPrompt: Design and Formulation
In this section, we first illustrate the overview of our proposed ConfusionPrompt, which consists of three critical components: decomposer, generator, and recomposer. Then, we develop a privacy model to formulate the requirements for privacy-preserving inference.
3.1 Overall Design
Denote $G_{s}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ as the LLM deployed as cloud service such as ChatGPT [45, 44, 8], where $\mathcal{V}^{*}$ is the vocabulary space. Instead of transmitting raw user prompt to the cloud, ConfusionPrompt introduces a privacy-preserving inference paradigm that significantly hinders third parties from inferring the original user prompt, even if they access the transmitted data. Our framework consists of six key steps:
Step 1: decomposition of the original prompt. We first introduce a decomposer $G_{d}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ that aims to decompose the user’s original prompt $p$ to a sequence of genuine sub-prompts $\boldsymbol{p}=[p_{1},p_{2},...,p_{|\boldsymbol{p}|}]$, resulting in fewer private attributes (i.e., private information) in each sub-prompt.
Step 2: generation of pseudo prompts. A pseudo prompt generator $G_{f}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ is introduced to generate a pseudo prompt given a genuine sub-prompt by replacing specified critical information. By generating multiple pseudo prompts for each genuine sub-prompt and mixing them together as a prompt group, we are able to hide the genuine sub-prompts, thus increasing the difficulty of being inferred. For the simplicity of privacy analysis, we design to enforce the pseudo prompt to be consistent with the genuine prompt in terms of syntactic structure.
Step 3: evaluation of privacy level. To ensure that the genuine prompt is safely hidden in the prompt group, we also need to design some criteria (e.g., semantic irrelevance) to evaluate the usability of generated pseudo-prompts. Based on these criteria, we iteratively sample and filter the pseudo-prompts until the prompt group meets the privacy requirement, which will be introduced in details in Section 3.2.
Step 4: communication with cloud server. The user transmits the prompt group to the server for LLM processing, and the server returns the corresponding response group. Throughout this process, the prompt group is designed to obscure the original user prompt, making it difficult for the server to infer the sensitives information, thereby safeguarding user’s privacy.
Step 5: retrieval of interested response. Since the user knows which sub-prompts are genuine, they can effortlessly extract the corresponding sub-responses while discarding those from pseudo-prompts.
Step 6: recomposition of sub-responses. Here, we introduce a recomposer $G_{r}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ that maps a sequence of sub-prompt-response pairs to a final response.
Figure 1: Overview of ConfusionPrompt.
3.2 Privacy Model
3.2.1 Rationale of Privacy Model
In this section, we define a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements that the group of pseudo-prompts should satisfy for privacy protection. This model provides guidance for evaluating privacy levels and training components (i.e., decomposer, generator, and recomposer). To explain the rationale of our privacy model, we follow [46] to quantify the privacy risk of queries exposed to the server.
Consider a set of prompts denoted as $\boldsymbol{P}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{n}\}$. For any $\boldsymbol{p}\in\boldsymbol{P}$, let $\pi(\boldsymbol{p})$ be the adversary’s prior probability that $\boldsymbol{p}$ is the genuine prompt provided by the user. Then the adversary’s posterior probability $\pi(\boldsymbol{p}|\boldsymbol{P})$ of the same event is:
$\pi(\boldsymbol{p}|\boldsymbol{P})=\frac{\pi(\boldsymbol{p})}{\sum_{\boldsymbol{p}^{\prime}\in\boldsymbol{P}}\pi(\boldsymbol{p}^{\prime})}.$ (1)
Let $\boldsymbol{p}^{0}$ denote the genuine prompt provided by the user. Then the privacy risk associated with the set of prompts $\boldsymbol{P}$ revealed to the server can be formulated as:
$\mathrm{Risk}(\boldsymbol{P})=\sum_{\boldsymbol{p}^{\prime}\in\boldsymbol{P}}\pi(\boldsymbol{p}^{\prime}|\boldsymbol{P})\cdot\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0}),$ (2)
where $\mathrm{Sim}$ is some measure of similarity between two prompts.
Equation 2 suggests three possible methods to minimize the privacy risk. First, reducing $\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0})$ for any $\boldsymbol{p}^{\prime}\neq\boldsymbol{p}^{0}$ leads to a decrease in $\mathrm{Risk}(\boldsymbol{P})$. Therefore, generating deceptive prompts that are semantically distinct from the genuine prompt can effectively lower the privacy risk. Second, $\mathrm{Risk}(\boldsymbol{P})$ is reduced when $\pi(\boldsymbol{p}^{\prime}|\boldsymbol{P})$ is high for prompts $\boldsymbol{p}^{\prime}$ that do not resemble $\boldsymbol{p}^{0}$. This implies that fabricated prompts should be crafted to appear as realistic as possible. Finally, $\mathrm{Risk}(\boldsymbol{P})$ decreases if we increase the number of prompts $\boldsymbol{p}^{\prime}$ with low $\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0})$. This can be achieved by increasing the number of fake prompts with different semantic meanings.
Based on the observations, we develop the following criteria for the fabricated prompts: (i) The fabricated attributes in the pseudo prompt should be semantically irrelevant to the original attributes in genuine prompt, so that the genuine user information can be effectively obfuscated [69]. (ii) The genuine prompt should be obfuscated by sufficient number of pseudo-prompts, and the attacker can not easily identify the real prompt from the combination pattern even with background knowledge. (iii) The pseudo-prompt should appear genuine rather than obviously fabricated. Thus it is crucial to maintain the fluency and reasonability of the pseudo-prompts.
3.2.2 Construction of Privacy Model
In the following, we first give the definitions of private attributes and their semantic relevance.
Definition 1 (Private Attributes).
Denote $\mathcal{U}$ as the attribute space, and $\mathcal{P}$ as the prompt space. Given a sub-prompt list $\boldsymbol{p}\in\mathcal{P}$, its private attributes are denoted as $U(\boldsymbol{p})=\{u_{1},u_{2},...,u_{m}\}$ for $m$ attributes.
Remark 2.
For conciseness, the prompt in Definition 1 could be a full prompt $p$, or a sequence of decomposed prompts with the same intention $\boldsymbol{p}=[p_{1},p_{2},...,p_{|\boldsymbol{p}|}]$.
The private attributes encompass a range of elements, including verbs, adjectives, and nouns, which extend beyond personally identifiable information (PIIs). Next, we define the semantic similarity between two private attributes.
Definition 3 (Attribute-attribute Similarity).
Given two attributes $u_{1}$, $u_{2}\in\mathcal{U}$, an attribute similarity function measures the semantic similarity between $u_{1}$ and $u_{2}$: $\text{Sim}(u_{1},u_{2}):\mathcal{U}\times\mathcal{U}\rightarrow\mathbb{R}$.
For a pair of genuine and pseudo-prompts, we define the correspondent attributes as follows:
Definition 4 (Correspondent Attributes).
Given two prompts $\boldsymbol{p}_{1}$, $\boldsymbol{p}_{2}\in\mathcal{P}$, the correspondent attribute of $u_{i}\in U(\boldsymbol{p}_{1})$ in $\boldsymbol{p}_{2}$, is defined as the attribute at the same syntactic position of $\boldsymbol{p}_{2}$, denoted by $\text{Corr}(u_{i},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$.
Remark 5.
The definition assumes that: (1) $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ have the same syntactic structure; (2) if $u_{i}$ occurs in $\boldsymbol{p}_{1}$ multiple times, the correspondent locations in $\boldsymbol{p}_{2}$ return the same attribute.
For example, given the original question ”What are the responsibilities of software engineers?” and private attribute ”software engineer”, the correspondent attribute in the fake question ”What are the responsibilities of school teachers?” would be ”school teacher”. Based on the correspondent attributes and attribute-attribute similarity, we can define the semantic similarity between two prompts below:
Definition 6 (Prompt-prompt Similarity).
Given two prompts $\boldsymbol{p}_{1}$, $\boldsymbol{p}_{2}\in\mathcal{P}$, their similarity can be defined by the similarity between each pair of correspondence attributes:
$\mathrm{Sim}(\boldsymbol{p}_{1},\boldsymbol{p}_{2})=\max_{u_{i}\in U(\boldsymbol{p}_{1})}\text{Sim}\left(u_{i},\text{Corr}(u_{i},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\right)$ (3)
Next, we define the significance of an attribute, which is tied to the likelihood of a curious server being able to identify the genuine attribute from a group of genuine and pseudo-prompts.
Definition 7 (Significance of Single Attribute).
Denote $\boldsymbol{P}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{n}\}$ as a group of prompts. The significance of an attribute $u\in U(\boldsymbol{p}_{i})$ related to a group of prompts $\boldsymbol{P}$ is defined as:
$\mathrm{Sig}(u,\boldsymbol{P})=\frac{\sum_{j=1}^{n}H(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j})}{n},$ (4)
where:
$H(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j})=\left\{\begin{array}[]{ll}1&\text{Corr}\left(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j}\right)=u\\ 0&\text{otherwise}\end{array}.\right$. (5)
In our setting, the group of prompts consists of one genuine prompt along with a collection of pseudo-prompts. The significance of a true attribute can be considered as the proportion of its occurrence within the group of prompts. In the following, we provide the definition of significance for multiple attributes.
Definition 8 (Significance of Attribute Set).
Suppose $\boldsymbol{p}_{0}$ denotes the genuine prompt. The significance of attribute set $U(\boldsymbol{p}_{i})$ related to a group of prompts $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ is defined as:
$\mathrm{Sig}(U(\boldsymbol{p}_{0}),\boldsymbol{P})$ (6)
[…]
${\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0h}),$
where:
$\bar{H}(\mathcal{V},p_{0h},p_{jh})=\left\{\begin{array}[]{ll}1&\mathrm{Corr}\left(u,p_{0h},p_{jh}\right)=u\ \forall u\in\mathcal{V}\\ &\mathrm{or}\ \mathcal{V}=\emptyset\\ 0&\mathrm{otherwise}\end{array},\right$. (7)
$p_{ih}$ is the $h$-th decomposed sub-prompt in the $i$-th prompt. In case that the attacker have prior knowledge of some private attributes, we introduce a notation $\mathcal{V}$ as a subset of private attributes in $p_{0h}$. Then we consider the proportion of a given attribute conditioned on any possible set of private attributes to bound the possibility of correctly identifying the target attribute with some background knowledge. Now, we provide the definition of genuineness for the third criterion:
Definition 9 (Genuineness).
Denote $D:\mathcal{P}\rightarrow\mathbb{R}$ as a function to discriminate between fabricated and genuine prompts. A larger value returned by $D$ indicates a higher likelihood of a prompt being genuine. The genuineness of a prompt $p\in\mathcal{P}$ can be defined as $\text{Genu}(p)=D(p)$.
Based on Definition 6, 8, and 9, we can formulate our proposed $(\lambda,\mu,\rho)$-privacy model as:
Definition 10 (User Privacy).
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are pseudo-prompts. If $\boldsymbol{P}$ satisfies the following requirements, then it is deemed that it can ensure $(\lambda,\mu,\rho)$-privacy of user prompt $\boldsymbol{p}_{0}$ with respect to discriminator $D$:
  • Each pseudo prompt should be semantically irrelavant to user prompt $\boldsymbol{p}_{0}$, i.e., $\forall i\in[1,n]$, $\text{Sim}(\boldsymbol{p}_{0},\boldsymbol{p}_{i})\leq\lambda$.
  • The users’ sensitive attributes should be obfuscated by sufficient pseudo-prompts, i.e., $\text{Sig}(U(\boldsymbol{p}_{0}),\boldsymbol{P})\leq\mu$.
  • Each pseudo-prompt should not be classified as fabricated prompt by discriminator $D$, i.e., $\forall i\in[1,n]$, $\text{Genu}(p_{i})=D(p_{i})\geq\rho$.
In .2, we provide the relationship between our privacy model and the upper bound on inference attacks.
3.2.3 Implication and Selection of Privacy Parameters
In this section, we provide a explanation for the three parameters in our privacy model:
  • Similarity parameter $\lambda$ controls the similarity between private attributes in genuine and pseudo-prompts. A lower $\lambda$ results in more dissimilar pseudo-prompts, thereby enhancing privacy protection. In our empirical analysis, we measure cosine similarity in the embedding space of private attributes. As shown in Figure 4, reducing $\lambda$ decreases the success rate of attribute inference attacks, with attack accuracy remaining sufficiently low for $\lambda<0.8$ under appropriate choices of the other two parameters (i.e., $\mu>10$ and $\rho=4$).
  • Significance parameter $\mu$ provides the bound on attack success rate of correctly identifying the target attribute. A lower $\mu$ necessitates generating a larger number of pseudo-prompts to obscure sensitive information, while providing higher level of privacy protection. The choice of $\mu$ depends on the user’s privacy requirements and the attribute space conditioned on the non-private information.
  • Genuineness parameter $\rho$ quantifies the degree of linguistic realism exhibited by individual synthetic prompts. In empirical analysis, we define the genuineness of a prompt on a four-point ordinal scale: 1 (incomprehensible), 2 (low quality), 3 (moderate fluency), and 4 (perfect fluency). We recommend the user to configure $\rho=4$, i.e., requiring each pseudo-prompt to satisfy perfect fluency, so that the attacker can not easily detect fabricated prompts.
4 ConfusionPrompt: User-Side Models
Following the overall design and the requirements in privacy model, this section introduces how to enable ConfusionPrompt through the design of user-side models, including decomposer, generator, and recomposer.
4.1 Decomposer
A decomposer is designed to decompose an original user prompt into several sub-prompts, such that the required number of pseudo-prompts to ensure the same level of privacy preservation can be significantly reduced. In the following, we theoretically demonstrate the benefit of a local decompose module in terms of complexity measured by the number of required pseudo-prompts.
4.1.1 Complexity Analysis
We first show the complexity for single-paragraph prompt, that is, the prompt is not decomposed.
Theorem 11 (Complexity for Single Paragraph).
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a single paragraph $|\boldsymbol{p}_{i}|=1$, $\forall i\in[1,n]$, meaning that each contains a single query. To achieve $(\lambda,\mu,\rho)$-privacy, it requires at least $n$ prompts:
$n\geq\left(\frac{1}{\mu}\right)^{|U(\boldsymbol{p}_{0})|},$ (8)
where $|U(\boldsymbol{p}_{0})|$ represents the number of private attributes in prompt $\boldsymbol{p}_{0}$.
Accordingly, we provide the following complexity analysis for decomposed prompts, which are a key component in our ConfusionPrompt.
Theorem 12 (Complexity for Decomposed Prompt).
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a sequence of decomposed prompts with the same intention $\boldsymbol{p}_{i}=[p_{i1},p_{i2},...,p_{il}]$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires at least $n$ prompts:
$n\geq\sum_{j=1}^{l}\left(\frac{1}{\mu}\right)^{|U(p_{0j})|},$ (9)
where $|U(p_{0j})|$ is the number of private attributes in the $j^{th}$ decomposed query, and $l$ denotes the number of decomposed queries in a prompt.
Specifically, if $|U(\boldsymbol{p}_{0})|=l$ and $|U(p_{0j})|=1$, $\forall j\in[1,l]$, meaning that each decomposed query contains a single attribute, then
$n\geq\frac{|U(\boldsymbol{p}_{0})|}{\mu}$ (10)
Theorem 11 and 12 compare the complexities between single prompt and decomposed sub-prompt, where the proof is defered to .1. In the ideal scenario, the complexity can be reduced to $\mathcal{O}\left(|U(\boldsymbol{p}_{0})|/\mu\right)$ after decomposition. An intuitive explanation is given in Figure 2, where a raw query containing three private attributes is decomposed into three sub-questions, each including a single attribute.
Figure 2: Real world example illustrating Theorem 11 and 12. The complexity is reduced from $(1/\mu)^{3}$ to $3/\mu$.
4.1.2 Requirements of Ideal Decomposer
Now, we can articulate the criteria for an ideal decomposer:
  • A fundamental requirement for the decomposition process is to follow the collectively exhaustive principle [53]. In other words, the collection of responses to decomposed sub-prompts should recover the whole response.
  • An ideal decomposer would optimize the complexity, i.e., the required number of pseudo-prompts, in two aspects: (i) each sub-prompt contains as few attributes as possible; (ii) each attribute appears in as few prompts as possible.
4.1.3 Model Training
We design a two-stage training approach to optimize the above objectives. In the first stage, we finetune a pretrained language model (LM) on demonstration data for decomposition task, where the LM is trained to generate decomposition given a full prompt. In the second stage, the LM is prompted to output multiple decompositions for each input. From the generations, a subset of preferred decompositions is selected to further finetune the model. The procedure to select preferred examples is as follows:
  • Evaluate whether each decomposition adheres to the collectively exhaustive principle, discarding any that do not meet this standard.
  • Among the qualified decompositions, assess the complexity of each using a predefined parameter $\mu$, and select the top $k$ examples for each prompt with the lowest complexity.
4.2 Generator
Generator takes an genuine prompt $p$ and its private attributes $U(p)$ as input and produces a pseudo prompt $p^{\prime}$ with replaced attributes. The generator will be run for multiple times to obtain a series of pseudo-prompts for each genuine prompt.
4.2.1 Requirements
The requirements for a pseudo-prompt suggest the following criteria for an ideal generator:
  • The generator should produce pseudo-prompt with fake information that is semantically irrelevant to the correspondent attributes in genuine prompt.
  • The pseudo-prompt should not be classified as fabricated prompt by a strong discriminator $D$.
4.2.2 Model Training
The generator is trained using a similar procedure described in Section 4.1.3. In the first stage, we finetune a pretrained LM using demonstration data for replacement task, where the LM is trained to generate new prompt with replaced attributes. In the second stage, we collect the preferred generations considering both semantic relevance and genuineness, with criteria given as follows:
  • Compute $s$, the similarity between the fabricated and original prompt.
  • Compute $f$, the genuineness score for the pseudo-prompt.
  • Calculate the overall score $\beta f-s$, where $\beta$ is the hyperparameter controlling the weight of the genuineness score.
  • Select the top $k$ examples for each prompt with the highest overall score.
4.3 Recomposer
The local re-composition LM combines the retrieved sub-prompts and sub-responses to produce the final whole response. We train a local recomposer on a collection of demonstration data with supervise learning. Specifically, given a complete prompt $p$ and a set of sub-prompt-response pairs $\left((p_{1},r_{1}),...(p_{k},r_{k})\right)$, the recomposer $G_{r}$ is trained to maximize $\log p_{G_{r}}\left(r|p,(p_{1},r_{1}),...(p_{k},r_{k})\right)$ where $r$ is the final response.
4.4 Algorithm
Algorithm 1 outlines the procedure for our private inference framework ConfusionPrompt. The client starts with decomposing the full prompt into sub-prompts and then generates a collection of qualified fake prompts that meet the privacy requirement. It can be derived immediately from the algorithm that our protocol meets $(\lambda,\mu,\rho)$-privacy.
Algorithm 1 ConfusionPrompt
1:Input: Original prompt $p$, privacy parameters $\lambda$, $\mu$, and $\rho$.
2:Output: Final response $r$.
3:User decomposes the original prompt into sub-tasks $\boldsymbol{p}_{0}=G_{d}(p)$.
4:for $p_{0h}\in\boldsymbol{p}_{0}$ do
5:  User keeps sampling from the generator until there are $n$ distinct pseudo prompts with $n$ given by 9, where each one $1\leq i\leq n$ has prompt similarity $\mathrm{Sim}(p_{0h},p_{ih})\leq\lambda$ and genuineness score $\mathrm{Genu}(p_{ih})\geq\rho$.
6:end for
7:User sends the group of genuine and pseudo prompts $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ to the server.
8:Server returns a collection of responses $(\boldsymbol{r}_{0},...,\boldsymbol{r}_{n})$.
9:User retrieves the response corresponding to the genuine prompt $\boldsymbol{r}_{0}$.
10:User obtains the final response $r$ using the local recomposition model.
In the following, we analyze the computational complexity for the user and server.
  • User computation: For $\boldsymbol{p}_{0}=[p_{01},p_{02},...,p_{0l}]$, the computation complexity can be broken as: (i) decompose the original prompt $O(l)$; (ii) generate pseudo-prompts $O\left(1/\alpha\sum_{j=1}^{l}\left(1/\mu\right)^{|U(p_{0j})|}\right)$, where $\alpha$ denotes the probability that a generated prompt satisfies the privacy requirement. In the ideal decomposition scenario, this reduces to $O\left(|U(\boldsymbol{p}_{0})|/(\alpha\mu)\right)$. (iii) recompose the responses $O(l)$. The primary computational overhead arises from pseudo-prompt generation.
  • Server computation: The server computation complexity is $O\left(\sum_{j=1}^{l}\left(1/\mu\right)^{|U(p_{0j})|}\right)$, which, in the ideal decomposition scenario, reduces to $O\left(|U(\boldsymbol{p}_{0})|/\mu\right)$. This directly corresponds to the user query cost.
5 Experiment
5.1 Experiment Setup
We evaluate our framework on three datasets: StrategyQA [22], MuSiQue [63], and MMLU [26]. These datasets encompass questions from a wide range of areas, such as business, medicine, and religion (see .3). We integrate ConfusionPrompt with different online LLMs including GPT-3.5-Turbo, GPT-4-Turbo, and GPT-4o [44]. We consider the privacy parameters $\lambda\in[0.5,0.8]$, $1/\mu\in[5,50]$, and $\rho\in[1,4]$ unless specified. Refer to .5 for the training of decomposer, generator, and recomposer.
We extract the private attributes of the three datasets through the steps below (see .4 for details):
  • Step 1: sample construction. We adopt a semi-automated method to extract private attributes from each sample. Specifically, we extract the entities using a combination of two NER methods: Spacy [58] and Flair [2]. Then we manually correct and supplement the private attributes for each query.
  • Step 2: model finetuning. An LLM is finetuned using the private attribute samples so that it outputs private attributes given a query.
  • Step 3: attribute extraction. The remaining queries are fed into the finetuned LLM to generate private attributes.
5.2 Empirical Privacy Evaluation
We simulate privacy attacks to investigate the privacy protection level under various combinations of privacy parameters. Existing privacy attacks can be categorized as: (i) membership inference attack that identifies whether a record is included in a model’s training dataset [57, 27], (ii) attribute inference attack that infers specific sensitive attributes [5, 4], and (iii) reconstruction attack that reconstructs the original data [74, 25].
The first attack is more associated with training time rather than inference time. Therefore, we focus on the latter two attacks, which are adapted to our ConfusionPrompt framework as follows:
Prompt Identification Attack: an attack that identifies the true query from a group of fake and true prompts. It is an adaptation of the reconstruction attack to our framework, where the attacker’s view is a set of prompts. Denote $\mathbf{q}=[\![q_{1}||q_{2}||...||q_{n}]\!]$ as the concatenated group of queries, and $k$ as the index for the true query. We finetune BART-large and Qwen3.5-4B classification models, denoted as $G_{\text{PIA}}$, to predict the index of the true query $\hat{k}=G_{\text{PIA}}(\mathbf{q})$.
Attribute Inference Attack: an attack that infers the sensitive features of records from either a group of prompts, or differentially privatized inputs. We rely on the twitter text dataset [65] to predict user gender based on the review.
Noted that prompt identification attack is specifically designed for our framework, and attribute inference attack can be applied to both ConfusionPrompt and DP-based methods.
5.3 Experiment Results
5.3.1 Privacy Experiments
Figure 3 visualizes the attack accuracy for prompt identification attack under various combinations of significance $\mu$ and genuineness $\rho$. It can be observed that: (a) Decreasing significance $\mu$ consistently reduces the attack accuracy. (b) As the genuineness threshold $\rho$ increases from 1 to 4, the attack accuracy approaches that of random guessing. The result reveals that it becomes harder for the attacker to distinguish the true query as the fake prompts become more realistic. In Figure 9, we investigate the impact of similarity threshold $\lambda$ on the attack success rate, and find no clear relationship between the two (see .7). Such privacy parameter could be more related to the attribute inference attack discussed later.
Figure 3: Prompt identification attack accuracy under various combinations of significance parameter $\mu$ and genuineness parameter $\rho$ for two attack models. Similarity parameter $\lambda$ is fixed at 0.5.
Figure 4 presents the attack accuracies of attribute inference attack for both ConfusionPrompt and LDP-based methods, from which we can make the following observations: (1) Paraphrase generally leads to higher attack accuracies compared to Text2text. This can be attributed to the fact that the semantic meanings of certain key words remain unchanged during the paraphrasing process. (2) The attack accuracies decrease to a plateau when $1/\mu\geq 15$ or $\epsilon\leq 1$. This suggests that beyond these thresholds, further increasing the privacy protection metrics does not significantly reduce the attack accuracy. (3) The attack accuracy for ConfusionPrompt with $1/\mu\geq 15$ is similar to that of Text2Text with $\epsilon\leq 1$, indicating certain alignment for the privacy protection between LDP-based methods and our framework. (4) Decreasing the similarity threshold $\lambda$ helps to reduce the attribute inference attack accuracies.
Figure 4: Attribute inference attack accuracy for ConfusionPrompt and LDP-based methods. Genuineness requirement $\rho$ is fixed at 4.
5.3.2 Utility Evaluation
For StrategyQA, we report the accuracy scores (ACC) and area under the roc curve (AUC) to assess the the classification task. For MuSiQue, we report the F1 score, ROUGE-L, and exact matching (EM) to evaluate the question answering task [70]. For MMLU, we report the ACC and F1 score to evaluate the multiple-choice task. We benchmark our framework with $\mu=15$, $\rho=4$, and $\lambda=0.5$ against three categories of baseline methods:
  • Local LLM. As ConfusionPrompt requires users to maintain multiple LMs for collaborative inference, it is essential to compare the performance of: (i) ConfusionPrompt utilizing online LLMs, and (ii) local inference using LLMs of comparable size to the LMs maintained on the user side. We thus evaluate the utility on two open-source LLMs: Llama2 [62] model with 7B parameters and Vicuna [11] model with 13B parameters.
  • LDP approaches. We consider two LDP approaches integrable with black-box LLMs: (i) Text-to-text privatization (Text2Text) [21], where the query is privatized with LDP by replacing each word with the perturbed token embeddings, and (ii) Paraphraser [64, 43] that paraphrases the prompt with LDP mechanism using a language model. The queries are privatized with LDP privacy budget $\epsilon=10$ before transmission to the server.
  • Direct Query that queries online GPTs with no privacy protection. The complete prompt is transmitted to the service provider without any confusion strategy.
Table 2: Utility comparisons between our proposed ConfusionPrompt and baselines.
Method Private StrategyQA MuSiQue MMLU
ACC AUC F1 roughL EM ACC F1
[…]
  The privacy parameters for ConfusionPrompt are fixed at $\mu=15$, $\rho=4$, and $\lambda=0.5$. The privacy budgets for LDP-based methods are set to $\epsilon=10$ and $\delta=10^{-4}$. Each value denotes the mean $\pm$ standard deviation under four rounds of experiments.
[…]
5.3.3 User Query Cost
Theorem 11 and 12 show that the decomposition module could reduce the number of pseudo-prompts required to be sent to the server. In this section, we empirically demonstrate practical benefits by analyzing the monetary cost of our ConfusionPrompt framework. In particular, we compute the monetary ratio, the ratio of the monetary cost for using ConfusionPrompt compared to the direct query method, both before and after the application of the decomposition module.
Figure 5 presents the query cost benefits achieved through our decomposition module. It can be observed that the complexity gains become increasingly pronounced as the level of privacy protection increases. This can be attributed to the fact that the complexity experiences polynomial growth in $1/\mu$ before decomposition, whereas it exhibits approximate linearity dependency on $1/\mu$ after decomposition. Our protocol offers substantial savings ranging from an average reduction of $3.3\times$ at $\mu=1/10$ to an average reduction of $18.2\times$ at $\mu=1/50$, demonstrating consistent trends in both online LLMs.
Figure 5: Monetary ratio of strategyQA, MuSiQue, and MMLU datasets for ConfusionPrompt and the framework without decomposition (ConfusionPrompt w/o Decomp) under various significance $\mu$. Decompostion in ConfusionPrompt substantially reduces the monetary cost, indicating its efficiency.
5.3.4 Computation Overhead Analysis
To validate the practicality of our method, we investigate the user-side computation overhead in terms of memory cost and sampling cost.
Memory Cost: In Figure 6, we present the memory cost required for the three local models, as well as that for local open-source LLMs. For MuSiQue, ConfusionPrompt (Base) utilizes BART-Base as the recomposer, while ConfusionPrompt (Large) employs BART-Large. For StrategyQA, ConfusionPrompt (Base) uses RoBERTa-Base as the recomposer, and ConfusionPrompt (Large) uses RoBERTa-Large.
From the results, we can see that our proposed ConfusionPrompt is significantly efficient compared to inferring large open-source language models on the user side. Even if using the large recomposer, it requires only $21.8\%$ and $41.8\%$ memory cost on average compared to run LLaMa2-7B and Vicuna 13B, respectively, on the user side. Additionally, ConfusionPrompt provides better utility than local LLM inference with lower memory requirement.
The empirical results also show the trade-off between memory requirement and utility for ConfusionPrompt. Using the base recomposer save the memory cost by an average of 17% compared with the large recomposer, with inferior performances. The suitable choice of local models depends on the user’s utility requirement and computational constraint.
Figure 6: User memory cost and utility on MuSiQue and StrategyQA for inference with 10 samples. ConfusionPrompt employs GPT-3.5-Turbo for online inference, with potential performance improvements using GPT-4-Turbo or GPT-4o.
Sampling Cost: Algorithm 1 suggests that the generator keeps generating fake prompts until the $(\lambda,\mu,\rho)$-privacy is satisfied. Therefore, it is crucial to consider the sampling time it takes for a generator to produce sufficient qualified pseudo-prompts. Figure 7 presents the number of generations to perform and the time cost for each generation under different levels of significance $\mu$. We consider three scenarios with varying combinations of genuineness parameter $\rho$ and similarity parameter $\lambda$.
For ConfusionPrompt, the sampling time scales approximately linear in significance requirement $1/\mu$. Our experiments indicate that under the most stringent privacy settings ($\rho=4$, $\lambda=0.5$), sampling is conducted an average of $2.5$ times per $1/\mu$. Relaxing the privacy constraints to $\rho=3$ and $\lambda=0.7$ reduces this ratio to approximately $2.1$. Note that $1/\mu$ can be treated as the minimal number of sampling times to achieve the privacy budget when the prompt contains only one private attribute.
Figure 7: Cost of generation in terms of time cost (seconds) for ConfusionPrompt and ConfusionPrompt w/o Decomp, under varying genuineness parameter $\rho$, similarity parameter $\lambda$, and significance parameter $\mu$.
Furthermore, we compare the generation cost for ConfusionPrompt and ConfusionPrompt w/o Decomp. We observe that decomposition reduces the sampling cost by an average of $5.88\times$ at $1/\mu=5$ and $26\times$ at $1/\mu=30$, as fewer pseudo-prompts are required to achieve the same privacy requirement.
5.3.5 Local Model Selection
In this section, we justify the selection of decomposer and recomposer. For decomposer, we obtain the BLEU between the generated decomposition and reference provided by the dataset. For recomposer, the models are evaluated in terms of the accuracy metrics described in Section 5.3.2, using the golden decomposition and sub-answer as training and testing inputs.
Table 3 presents the experiment results on MuSiQue using models from three categories (i.e., BART, T5, and Flan-T5) of varying sizes. While Flan-T5-large exhibits the best performance for both tasks, we select BART-large as the decomposition and recomposition model due to its balanced trade-off between model size and performance.
Table 3: Parameter size and performance on both decomposition and recomposition tasks across various models for MuSiQue.
Decomposer Recomposer
BLEU F1 RougeL EM
BART-base (139M) 0.314 0.563 0.564 0.446
BART-large (406M) 0.402 0.640 0.641 0.525
T5-base (223M) 0.343 0.581 0.580 0.484
T5-large (737M) 0.408 0.646 0.645 0.554
Flan-T5-base (248M) 0.367 0.608 0.609 0.518
Flan-T5-large (783M) 0.458 0.671 0.671 0.577
6 Discussion
ConfusionPrompt focuses on the privacy protection for textual data during privacy-preserving LLM inference. In this section, we provide further clarifications and discussions on practical issues of our framework.
6.1 Protection Capability of Privacy Model
Our $(\lambda,\mu,\rho)$-privacy model establishes requirements for fabricated prompts to mitigate privacy risks. Rather than offering a worst-case privacy guarantee such as LDP, the model quantifies the criteria that a group of prompts should satisfy. Our theoretical and empirical experiments associate the privacy parameters with attack success rate. It is important to note that the privacy model should be regarded as an enhancement over the insecure status quo, rather than a substitute for formal privacy guarantees.
Alternative privacy models could be explored to enhance privacy protection in future studies. One potential direction is to analyze the distribution of sensitive attributes conditioned on the non-sensitive components, ensuring that the set of pseudo-prompts closely approximates the ground-truth distribution.
6.2 Trade-off between Privacy and Query Cost
Whereas traditional perturbation-based methods trade privacy for utility, our ConfusionPrompt trades privacy for query complexity while providing accurate responses. This approach is particularly suitable for clients with high privacy demands who can accommodate these costs, such as finance and healthcare organizations. Additionally, although query costs rise linearly with the number of tokens in API services, users of chatbot services can submit unlimited requests under a fixed subscription fee, allowing for enhanced privacy protection at a manageable cost.
The query complexity could be further reduced by considering more condensed representation of the batch of pseudo-prompt (see .8). While the complexity still remains at least linearly dependent on the significance requirement $1/\mu$, future research may explore approaches to achieve a sub-linear dependence on $1/\mu$.
6.3 Generation of Pseudo-Prompts
The privacy protection of our framework relies on the quality of pseudo-prompts in effectively concealing sensitive attributes. In this work, we define the quality of fabricated prompts based on three key privacy considerations: the significance of attributes, the genuineness of the pseudo-query, and its similarity to the genuine prompt. Additionally, the generated prompts should maintain a consistent semantic structure for the non-sensitive components. The privacy requirement and assumption are satisfied by selecting filtered prompts from the generated candidates. Our approach employs a lightweight local generator to produce fabricated prompts; however, this may be insufficient, particularly when generating a large number of pseudo-prompts under stringent privacy requirements. To address this, we propose an alternative method in .9 that leverages a powerful online LLM for pseudo-prompt generation while ensuring that private attributes remain hidden from the LLM owner.
In Section 5.3.1, we present results from prompt identification attack, demonstrating the difficulty in distinguishing real prompt from pseudo-prompts under appropriate choice of privacy parameters. While more advanced discriminators may achieve higher attack success rates, developing stronger attackers is left for future work. This may involve using LLMs as discriminators and integrating domain knowledge. To further enhance privacy protection, future improvements could: (1) enhance quality assessment by incorporating additional privacy requirement, such as the diversity of pseudo-prompts, to strengthen privacy protection; (2) refine the evaluation of genuineness and similarity in experiments by considering metrics beyond fluency scores and embedding cosine similarity; (3) train the generator using adversarial training techniques and employ a set of sophisticated discriminators to select qualified pseudo-prompts, thereby improving the robustness of pseudo-prompt generation.
6.4 User Control over Sensitive Data
For privacy protection, we implement the decomposer and generator locally on the user side, granting users complete control over their sensitive data. This setup allows users to customize pseudo-prompt generation based on their privacy and ethical requirements, domain expertise, and computational resources. In particular:
  • Users can freely set the privacy parameters $(\lambda,\mu,\rho)$. For stringent privacy needs, users might: (1) assign a lower value to the significance parameter $\mu$, resulting in a larger set of pseudo-prompts; (2) set a lower similarity parameter $\lambda$ and a higher genuineness parameter $\rho$ by adjusting the filtering threshold.
  • During pseudo-prompt generation, users can filter prompts to achieve the requirement of $\lambda$ and $\rho$ through: (1) utilizing a local similarity and discrimination model (see Appendix .6), ensuring intermediate results remain on the user’s device for enhanced privacy; (2) employing another remote LLM to assess pseudo-prompt quality, offloading computational tasks to a server, which maintains privacy if the two LLM servers do not collaborate. Additionally, users can manually review and modify prompt groups before sending them to the server.
7 Conclusion
This paper proposes a private inference framework for online LLMs, termed ConfusionPrompt. We deploy three local models on the user side: (i) decomposer that maps an original prompt to a sequence of sub-prompts, (ii) generator that produces pseudo-prompts by replacing private attributes in the genuine sub-prompts, (iii) recomposer that maps the decomposed prompt-response pairs from the cloud service to the final response. Such design endows our framework with advantages over previous protocols that: (i) it can be seamlessly integrated with existing black-box LLMs where the service providers have no need to modify their already-built framework, and (ii) it achieves better privacy-utility trade-off than existing LDP-based privatization methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirement for a privacy-preserving group of prompts, and accordingly provide a complexity analysis to demonstrate the benefits of the decomposition module. Experiments show that our ConfusionPrompt outperforms local inference with open-source models, while reducing memory consumption by at least 62% compared to open-source LLMs. Furthermore, it achieves over 42% higher utility than LDP-based methods.